OpenStudy (anonymous):

Find the domain of the exponential function f(x) = (1/3)x. A. Any real numbers B. Positive real numbers C. Negative real numbers D. Non-zero real numbers Find the domain of the exponential function f(x) = -2(3)x. A. Any real numbers B. Positive real numbers C. Negative real numbers D. Non-zero real numbers

OpenStudy (anonymous):

ummm so are the two functions supposed to be e^.3333x and e^-6x ???

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

well the exponential function is y = e^x... so I am not sure exactly what those are

OpenStudy (nubeer):

any real numbers.

OpenStudy (anonymous):

@dasDaName Once again, these aren't exponential functions! Are you sure you wrote down the questions correctly? Maybe you forgot to write exponents...

OpenStudy (anonymous):

it might be (1/3)^x

OpenStudy (anonymous):

Maybe.

OpenStudy (anonymous):

yes technofable,

OpenStudy (anonymous):

i thought writing an X without the '^' power was okay, nvm its (1/3)^x

OpenStudy (anonymous):

s its Find the domain of the exponential function f(x) = (1/3)^x. Find the domain of the exponential function f(x) = -2(3)^x

OpenStudy (anonymous):

first one should be all positive numbers. u could just plug any number in the exponent and verify it. i'm not sure what the second one says though.

OpenStudy (anonymous):

is it \[2\times3^x\]?

OpenStudy (anonymous):

second one yes, but -2*3^x

OpenStudy (anonymous):

then it should be all negative real

OpenStudy (anonymous):

thanks technfable